OpenStudy (anonymous):

what is the slope intercept form for the following equation: -2x-11y=5

OpenStudy (anonymous):

@WildQueen @a1234

OpenStudy (tkhunny):

This requires solving for y. Go!

OpenStudy (astrophysics):

Slope intercept form is y = mx+b

OpenStudy (astrophysics):

So yes, solve for y as tk mentioned, m here is your slope, and b is the y - intercept

OpenStudy (anonymous):

wouldnt the answer be y=-2/11x-5/11 for the equation and slope:-2/11 and y intercept -5/11???

OpenStudy (astrophysics):

Yes :)

OpenStudy (anonymous):

:))) do you think you can help me on a couple more?

OpenStudy (anonymous):

@Astrophysics

OpenStudy (anonymous):

@alexandra_annabella @tkhunny @imqwerty

OpenStudy (anonymous):

which graph best represents the equation -5x+29=2?

OpenStudy (tkhunny):

Are you sure you wrote it correctly? Why not change to slope-intercept form and see for yourself?

imqwerty (imqwerty):

in this equation nothing is given about y so y can be anything but x is constant can u solve this equation to find out the value of x?? :)

OpenStudy (astrophysics):

Probably meant -5x+29y=2?

OpenStudy (anonymous):

no im having trouble with a quiz that i have taken 3 times already and this is the last time i can take it and pass

imqwerty (imqwerty):

find out the value of x and it will be like this-